# The Fréchet Distance Unleashed: Approximating a Dog with a Frog

### v3

@@ 6 Conclusions and future work
In this work, we have demonstrated that our variant of the Fréchet distance is both theoretically efficient and viable in practice. We provided standard libraries for the algorithms in Julia and Python. Furthermore, our libraries compute the retractable Fréchet distance/morphing, which seems to perform quite well in practice, and is better at handling noise than the original Fréchet distance. This suggests that our algorithm may be useful in applications where the runtime is dominated by Fréchet distance computations. More generally, we believe the high quality of our implementations, and their easy availability via standard libraries in Julia and Python should make our code and algorithms easily accessible for casual users.
@@
Using the implementation given by Bringmann et al. bkn-wdfpa-21, there has been substantial empirical work on efficient clustering of curves. Notably, Buchin et al. bdvn-kcbct-19 empirically study algorithms for $(k,\ell)$-center clustering. We omit the formal definition of this problem here, but of note is the large number of Fréchet distance computations required to determine the furthest curve from the set of existing curves chosen as centers. We believe that our algorithm is very well suited for this task, as our heap-based implementation can be easily modified to stop early if the Fréchet distance goes beyond a given threshold.
Similarly, there has been prior empirical work on $(k,\ell)$-median clustering bbknp-klmct-20 under the dynamic time warping distance, and we believe our algorithm would perform well in this application as well.
@@
A strategy that works quite well for the task of Section 2.2.2 is to precompute, for all input curves, the fast simplification extractor of Lemma 5.25. Then, given a query, one can simplify the two curves quickly to the right resolution $\delta$ (initially guessed to be some large value) using this extractor, then apply the algorithm of Lemma 5.27 to get a simplification with few vertices. With these simplifications, then compute the Fréchet distance, and repeat the process, with smaller $\delta$, till the query is resolved. By caching the simplifications of the input curves, one can achieve better speedup if there are multiple computations involving the same curve.
@@ Appendix A The elevation function
We need some standard properties of the elevation function. We prove some of them here, so that our presentation would be self contained.

### v4

@@ 6 Conclusions and future work
In this work, we have demonstrated that our variant of the Fréchet distance is both theoretically efficient and viable in practice. We provided standard libraries for the algorithms in Julia and Python. Furthermore, our libraries compute the retractable Fréchet distance/morphing, which seems to perform quite well in practice, and is better at handling noise than the original Fréchet distance. This suggests that our algorithm may be helpful in applications where Fréchet distance computations dominate the runtime, especially when approximation is acceptable. More generally, we believe the high quality of our implementations, and their easy availability via standard libraries in Julia and Python, should make our code and algorithms easily accessible for casual users.
@@
Using the implementation given by Bringmann et al. [BKN21], there has been substantial empirical work on efficient clustering of curves. Notably, Buchin et al. [BDLN19] empirically study algorithms for $(k,\ell)$-center clustering. We omit the formal definition of this problem here, but of note is the large number of Fréchet distance computations required to determine the furthest curve from the set of existing curves chosen as centers. We believe that our algorithm is very well suited for this task, as our heap-based implementation can be easily modified to stop early if the Fréchet distance goes beyond a given threshold.
Similarly, there has been prior empirical work on $(k,\ell)$-median clustering [BBK+20] under the dynamic time warping distance, and we believe our algorithm would perform well in this application as well.
@@
A strategy that works quite well for working on a large database of curves is to precompute, for all input curves, the fast simplification extractor of Lemma 5.25. Then, given a query, one can simplify the two curves quickly to the correct resolution $\delta$ (initially guessed to be some large value) using this extractor, then apply the algorithm of Lemma 5.27 to get a simplification with few vertices. With these simplifications, then compute the Fréchet distance, and repeat the process, with smaller $\delta$, till the query is resolved. By caching the simplifications of the input curves, one can achieve better speedup if there are multiple computations involving the same curve.
When our algorithm performs badly?
Our algorithm performs poorly when many refinement rounds are required, and even worse when there are many different paths that are more or less equivalent, and the refinement causes the morphing to jump around the free space during refinement. Specifically, such bad examples are made out of “one-dimensional” zigzags where the distance is large, just because one of the curves has one more zigzag than the other, or even if the zigzags match but are unfavorable. Such an example also implies that many rounds of refinements are required before the algorithm gets even remotely close to the optimal Fréchet distance. See Figure 6.1. Of course, such inputs are not in any way real-world input, and represent the gap between the theory and practice of computing the Fréchet distance.
Figure 6.1: When the new algorithm does badly – the implementation required about 120 refinement rounds before converging, as the morphing kept leaping around the free space diagram. See here for more details and animations. This example can be made worse by scaling the $y$-axis further down.
@@ Appendix A The elevation function
We need some standard properties of the elevation function. We prove some of them here, so that our presentation would be self-contained.